OpenStudy (anonymous):

an investment account started at 3500 and collected annually compounded interest fro 1980 until 2005. At the end of 2005 the account was worth 78000. write an equation for the value of th investment, t years after1980

OpenStudy (anonymous):

first you need the rate of interest. solve \[3500(1+r)^{25}=78000\] for \(r\) . do you know who to do that?

OpenStudy (anonymous):

no i dont know

OpenStudy (anonymous):

1) divide by 3500 first \[(1+r)^{20}=\frac{156}{7}\] 2) take the 20th root of both sides \[1+r=\sqrt[20]{\frac{156}{7}}=1.1679\] 3) subtract 1 get \[r=.1679\] that seems like an awfully high rate of interest. let me check again

OpenStudy (anonymous):

yeah i made a mistake, it should be \[1+r=\sqrt[25]{\frac{156}{7}}=1.132\] and so \[r=.132\] still very hight but it think this one is right

OpenStudy (anonymous):

thanks